# Curriculum Negative Mining For Temporal Networks

### v1

@@ 5.3 Experimental Results
The inductive setting allows us to compare our method’s generalizability against existing methods by evaluating their performances on unseen nodes. When comparing inductive to transductive settings, the average ranks of ENS and recent sampling worsen, suggesting that their heuristic strategies might not be generalizable and could lead to overfitting on seen nodes. On the contrary, the average rank of random sampling improves, indicating its robustness despite relatively weaker performance. Finally, our CurNM method consistently achieves the best across 9 datasets and delivers a higher average rank in the inductive setting compared to the transductive setting. This confirms that our method not only performs well but is also robust and highly generalizable.
At the dataset level, our method consistently outperforms or matches the existing methods in almost all datasets, with insignificant declines in only two cases (again, with myket being an outlier).
@@ 5.4 Complexity Analysis
Theoretically, CurNM adds a negative mining process with a time complexity of $O(|E|\times n\times d)$ to TGNNs, which originally have a complexity of $O(L\times|E|\times n\times d)$. Here, $L$ denotes the number of layers, $|E|$ is the number of edges, and $n$ represents the number of positive and negative samples per node, with nonlinear operations simplified to $O(d)$.
Empirically, CurNM maintains high efficiency by limiting $n$ with a relatively small negative pool, sized at $M=8$. Table 3 presents a runtime comparison between CurNM and baseline methods on the TGN model, which demonstrates the best overall performance among the three TGNNs evaluated. According to Table 3, CurNM adds minimal computational load compared to basic random and recent sampling, consistently requiring about twice the time regardless of dataset size. Moreover, it is significantly more efficient than the state-of-the-art ENS on large datasets, requiring approximately one-seventh of the time on Flights, the largest dataset tested.
Table 3: Comparison of average runtimes per epoch between CurNM and baseline methods on the TGN model. $Ratio_{Rand}$, $Ratio_{Rec}$, and $Ratio_{ENS}$ represent the ratio of CurNM runtime to Random, Recent, and ENS runtimes, respectively.
Dataset Random $Ratio_{Rand}$ Recent $Ratio_{Rec}$ ENS $Ratio_{ENS}$ CurNM
uci 8.0 1.86 8.6 1.73 10.8 1.38 14.9
USLegis 8.8 1.86 8.6 1.91 12.1 1.36 16.4
CanParl 11.3 1.47 11.4 1.46 17.1 0.97 16.6
enron 16.4 1.7 17.0 1.64 17.7 1.58 27.9
WIKI 22.7 1.45 23.4 1.41 28.5 1.15 32.9
MOOC 39.9 2.09 43.4 1.92 152.7 0.55 83.3
UNtrade 65.6 1.83 71.8 1.67 154.6 0.78 120.1
REDDIT 84.9 1.76 116.7 1.28 290.8 0.51 149.0
myket 79.3 2.18 89.2 1.94 1133.9 0.15 172.8
UNvote 116.0 1.91 138.4 1.6 422.3 0.52 221.0
LASTFM 161.2 2.29 156.2 2.36 1019.0 0.36 368.8
Flights 221.8 2.1 231.7 2.01 3222.5 0.14 466.6
@@ 5.5 Ablation Study
The ablation study detailed in Figure 2 examines the effects of adaptive $\pi_{e}$, annealing random negatives, hard negative cache, factor disentanglement, and temporal-aware embeddings in our method on the TGN model. Due to space constraints, we present our study using six datasets here. In both transductive and inductive settings, the complete CurNM model delivers the most robust performance across the datasets. The most significant decline in performance is observed when random negatives are removed, validating our claim that a steep difficulty curve can confuse the model. In datasets like USLegis, CanParl, and enron, where our model achieves strong results, the removal of temporal-aware embeddings and adaptive $\pi_{e}$ also leads to substantial performance drops. This underscores their critical role in addressing the networks’ positive sparsity and positive shift. Although the impact of removing disentanglement and hard negative cache is less pronounced, their absence can still result in up to a 3% performance decline, suggesting their importance in enhancing the granularity of learning within the model.
Figure 2: Ablation studies on different components.
@@ 5.6 Hyper-Parameter Sensitivity
We investigate our assumptions about the structures of $\alpha_{e},\beta_{e}$, and $\delta_{e}$ by testing an increasing strategy (e.g., $\alpha_{e}=\min(\frac{e}{E^{\prime}},1)$), a decreasing strategy (e.g., $\alpha_{e}=\max(0,1-\frac{e}{E^{\prime}})$), and a constant strategy (e.g., $\alpha_{e}=0.5$). Our findings in Figure 3 confirm that, as expected, our current approach yields the most robust results. This underscores the effectiveness of our dynamic strategies, which progressively guide the model to focus on relevant factors of informative nodes, thereby achieving a smoother and more advantageous learning curve.
@@
Under the framework where $\pi_{e}$ should shrink as learning progresses, we explore two approaches: a linear strategy and an adaptive strategy. The linear strategy consistently reduces the $\pi_{e}$ by a set amount each epoch, whereas the adaptive strategy only does so when there’s an improvement in the selected performance metric. For our baseline, both methods start with $\pi_{e}=100\%$, with a shrinkage rate $\gamma_{\text{shrink}}=3\%$, and the $\pi_{e}$ can decrease to a minimum of $10\%$. The choice of $\gamma_{\text{shrink}}=3\%$ is based on that models generally stabilize after approximately $30$ epochs, and having a larger $\pi_{e}$ before this phase to capture more information would be beneficial. The adaptive strategy outperforms the linear one. We then experiment with a slower ($\gamma_{\text{shrink}}=2\%$) and a faster shrinkage rate ($\gamma_{\text{shrink}}=5\%$). Figure 3 shows that $\gamma_{\text{shrink}}=3\%$ yields the best results.
Figure 3: Parameter sensitivity studies on $\alpha_{e},\beta_{e},\delta_{e},\text{ and }\pi_{e}$.
@@ 6 Conclusion
In this paper, we identify two unique challenges of negative sampling in temporal networks – positive sparsity and positive shift. Our proposed CurNM method employs a dynamically updated negative pool and a curriculum learning strategy to adaptively adjust the difficulty level of negative samples during training, thus addressing the common challenge of positive sparsity. Furthermore, we implement a temporal-aware negative selection function to effectively capture positive shift. Our method’s efficacy is demonstrated through extensive experiments against 3 algorithms and 12 datasets. Although CurNM demonstrates superior performance, it requires tuning several hyper-parameters to optimize its performance. In the future, we aim to simplify these designs to reduce the need for extensive parameter tuning.
@@ Acknowledgements
This work was supported by the Joint Funds of the Zhejiang Provincial Natural Science Foundation of China [grant number LHZSD24F020001]; Zhejiang Province ”JianBingLingYan+X” Research and Development Plan [grant number 2024C01114]; Ningbo Natural Science Foundation [grant number 2023J281]; and Zhejiang Province High-Level Talents Special Support Program ”Leading Talent of Technological Innovation of Ten-Thousands Talents Program” [grant number 2022R52046].

### v2

@@ 5.4.2 Inductive Setting
The inductive setting allows us to compare our method’s generalizability against existing methods by evaluating their performances on unseen nodes. When comparing inductive to transductive settings, the average ranks of ENS and recent sampling worsen, suggesting that their heuristic strategies might not be generalizable and could lead to overfitting on seen nodes. On the contrary, the average rank of random sampling improves, indicating its robustness despite relatively weaker performance. CurNM also achieves a better average rank in the inductive setting, confirming that it is not only effective but also robust and highly generalizable.
At the model level
According to Table 4, CurNM achieves statistically significant average improvements when combined with TGN across all datasets, with an average gain of 6.6%. It also shows no significant deterioration when used with JODIE, yielding an average gain of 3.9%. DySAT remains slightly weaker, exhibiting similar patterns as in the transductive setting.
At the dataset level
CurNM also follows similar overall trends. The main difference is the slight performance drop in CanParl and MOOC, due to the relatively stronger performance of DySAT - an architecture less compatible with CurNM’s design. Nevertheless, even in these two cases, CurNM falls short of the best method by no more than 2%.
Table 4: Average Precision of link prediction across 12 datasets and three TGNNs in the inductive setting. “AvgImp” denotes the average percentage improvement of CurNS compared to other negative sampling methods. Cells highlighted in red indicate that CurNS achieved a statistically significant improvement, while blue cells indicate a significant deterioration. Uncolored cells reflect no significant difference.
NaN indicates failed results. “Avg Rank” refers to the average ranks of each sampling method. Dataset Model Random Recent ENS TASER CurNM AvgImp (%) uci DySAT 0.726 0.727 0.737 0.704 0.748 3.335 JODIE 0.636 0.542 0.648 0.590 0.652 8.403 TGN 0.677 0.662 0.700 0.667 0.705 4.351 USLegis DySAT 0.480 0.584 0.510 0.503 0.516 0.041 JODIE 0.597 0.596 0.618 0.664 0.602 -2.575 TGN 0.763 0.730 0.789 0.691 0.817 10.185 CanParl DySAT 0.672 0.630 0.674 0.667 0.662 0.193 JODIE 0.533 0.539 0.530 0.660 0.564 0.645 TGN 0.626 0.624 0.604 0.610 0.666 8.228 enron DySAT 0.704 0.698 0.703 0.693 0.700 0.065 JODIE 0.724 0.693 0.716 0.543 0.725 9.888 TGN 0.730 0.811 0.769 0.709 0.827 9.809 WIKI DySAT 0.821 0.804 0.823 0.813 0.822 0.826 JODIE 0.701 0.701 0.712 0.643 0.698 1.426 TGN 0.856 0.837 0.860 0.848 0.875 2.914 MOOC DySAT 0.684 0.592 0.687 0.710 0.653 -1.758 JODIE 0.580 0.574 0.576 0.545 0.587 3.225 TGN 0.688 0.653 0.665 0.652 0.698 5.155 UNtrade DySAT 0.538 0.556 0.539 0.511 0.545 1.674 JODIE 0.609 0.680 0.613 0.606 0.623 -0.415 TGN 0.661 0.665 0.660 0.651 0.683 3.565 REDDIT DySAT 0.823 0.814 0.830 0.828 0.817 -0.813 JODIE 0.787 0.777 0.796 0.814 0.810 2.149 TGN 0.874 0.853 0.872 0.880 0.880 1.247 myket DySAT 0.596 0.689 0.592 0.597 0.597 -3.024 JODIE 0.530 0.533 0.550 0.551 0.531 -1.857 TGN 0.582 0.539 0.578 0.577 0.585 2.847 UNvote DySAT 0.511 0.499 0.509 0.513 0.519 2.137 JODIE 0.694 0.635 0.698 0.521 0.689 9.604 TGN 0.650 0.594 0.660 0.506 0.672 12.747 LASTFM DySAT 0.709 0.622 0.709 0.705 0.748 9.325 JODIE 0.549 0.526 0.557 0.564 0.581 5.884 TGN 0.706 0.596 0.702 0.601 0.690 6.663 Flights DySAT 0.710 0.689 nan 0.701 0.507 -27.584 JODIE 0.656 0.605 nan 0.662 0.705 10.112 TGN 0.696 0.665 nan 0.669 0.751 10.988 Avg Rank 2.94 3.81 2.61 3.53 1.92
@@ 5.5 Complexity Analysis
Theoretically, CurNM introduces a negative mining process with a time complexity of $O(|E|\times(1+M)\times d)$ to TGNNs, which originally have a complexity of $O(L\times|E|\times n\times d)$. Here, $L$ denotes the number of layers, $|E|$ is the number of edges, and $(1+M)$ and $n$ represent the total number of positive plus negative samples per edge with and without CurNM, respectively. Nonlinear operations are simplified to $O(d)$.
Empirically, CurNM maintains high efficiency by limiting $M$ to a relatively small negative pool size. Table 5 presents a runtime comparison between CurNM and baseline methods on the TGN model, which demonstrates the best overall performance among the three TGNNs evaluated. We first fix $M=8$ when comparing CurNM with the baselines  [^3] Based on our grid search, we found that a pool size of $M=8$ yields the best results., and later examine how varying $M$ affects runtime.
According to Table 5, CurNM adds minimal computational load compared to random sampling, recent sampling, and TASER, consistently requiring about twice the time regardless of dataset size. Moreover, it is significantly more efficient than ENS on large datasets, requiring approximately one-tenth of the time on Flights, the largest dataset tested.
Table 5: Comparison of average runtimes per epoch between CurNM and baseline methods on the TGN model. “Ratio” represents the ratio of CurNM runtime to each baseline method.
Dataset Random Ratio Recent Ratio ENS Ratio TASER Ratio CurNM
uci 7.7 1.6 8.1 1.6 9.5 1.3 7.9 1.6 13.2
USLegis 7.6 2.1 8.2 1.9 10.7 1.5 7.9 2.0 13.0
CanParl 8.2 2.4 8.8 2.2 16.1 1.2 8.3 2.3 15.7
enron 13.0 2.2 14.0 2.1 16.5 1.8 12.8 2.3 23.8
WIKI 18.7 1.8 20.1 1.7 27.5 1.2 18.6 1.8 29.1
MOOC 42.1 2.1 47.4 1.9 169.2 0.5 42.7 2.1 73.0
UNtrade 54.7 2.3 63.6 2.0 148.3 0.8 58.4 2.1 108.0
REDDIT 77.0 2.1 86.8 1.9 310.7 0.5 78.9 2.0 144.1
myket 79.8 2.3 93.7 2.0 1156.6 0.2 83.3 2.2 158.9
UNvote 119.1 2.0 134.1 1.8 350.1 0.7 123.2 1.9 205.3
LASTFM 175.1 1.8 178.1 1.7 987.4 0.3 171.7 1.8 333.1
Flights 224.3 1.8 223.1 1.8 3236.7 0.1 219.8 1.9 446.2
According to Table 6, increasing $M$ has little impact on the runtime. This is primarily because the negative pool size only affects the simple prediction layer, rather than the more computationally expensive graph embedding process, and thus contributes minimally to the overall model complexity.
Table 6: Comparison of average runtime per epoch with different negative pool sizes.
Dataset M=4 M=8 M=12 M=16
uci 13.6 13.2 13.6 15.0
USLegis 14.1 13.0 13.0 16.2
CanParl 15.9 15.7 16.1 16.3
enron 26.0 23.8 24.4 27.3
WIKI 33.6 29.1 29.7 36.3
MOOC 85.2 73.0 78.2 85.6
UNtrade 111.2 108.0 111.1 109.2
REDDIT 149.9 144.1 158.0 179.7
myket 152.0 158.9 177.3 183.3
UNvote 189.3 205.3 209.1 224.6
LASTFM 289.3 333.1 335.4 363.1
Flights 427.1 446.2 461.9 510.9
@@ 5.6 Ablation Study
The ablation study detailed in Figure 2 examines the effects of adaptive $\pi_{e}$, annealing random negatives, hard negative cache, factor disentanglement, and temporal-aware embeddings in our method on the TGN model. Due to space constraints, we present our study using eight datasets here. In both transductive and inductive settings, the complete CurNM model delivers the most robust performance across the datasets.
The most significant decline in performance is observed when random negatives are removed, validating our claim that a steep difficulty curve can confuse the model. In datasets like USLegis, CanParl, and enron, where our model achieves strong results, the absence of temporal-aware embeddings and adaptive $\pi_{e}$ also leads to substantial performance drops. This underscores their critical role in addressing the networks’ positive sparsity and positive shift. Although the impact of removing disentanglement and hard negative cache is less pronounced, their absence can still result in up to a 9% performance decline, suggesting their importance in enhancing the granularity of learning within the model.
Figure 2: Ablation studies on different components presented as percentage differences relative to the complete CurNM design.
@@ 5.7 Hyper-Parameter Sensitivity
This section tests our hypotheses underlying the key parameters that govern the curriculum learning process.
@@
We investigate our assumptions about the structures of $\alpha_{e},\beta_{e}$, and $\delta_{e}$ by testing an increasing strategy (e.g., $\alpha_{e}=\min(\frac{e}{E^{\prime}},1)$), a decreasing strategy (e.g., $\alpha_{e}=\max(0,1-\frac{e}{E^{\prime}})$), and a constant strategy (e.g., $\alpha_{e}=0.5$). Our findings in Figure 3 confirm that, as expected, our proposed approach yields the most robust results. This highlights the effectiveness of our dynamic strategies in guiding the model gradually toward more relevant and difficult nodes, leading to a smoother and more efficient learning curve.
@@
Under the framework where $\pi_{e}$ should shrink as learning progresses, we explore two approaches: a linear strategy and an adaptive strategy. The linear strategy consistently reduces the $\pi_{e}$ by a set amount each epoch, whereas the adaptive strategy only does so when there’s an improvement in the selected performance metric. For our baseline, both methods start with $\pi_{e}=100\%$, with a shrinkage factor $\gamma_{\text{shrink}}=3\%$, and the $\pi_{e}$ can decrease to a minimum of $10\%$. The choice of $\gamma_{\text{shrink}}=3\%$ is based on that models generally stabilize after approximately $30$ epochs, and having a larger $\pi_{e}$ before this phase to capture more information would be beneficial. The adaptive strategy outperforms the linear one. We then experiment with a slower ($\gamma_{\text{shrink}}=2\%$) and a faster shrinkage factor ($\gamma_{\text{shrink}}=5\%$). Figure 3 shows that $\gamma_{\text{shrink}}=3\%$ yields the best results.
Figure 3: Parameter sensitivity studies on $\alpha_{e}$, $\beta_{e}$, $\delta_{e}$, and $\pi_{e}$ presented as percentage differences relative to the complete CurNM design.
@@ 6 Conclusion
In this paper, we identify two unique challenges of negative sampling in temporal networks – positive sparsity and positive shift. Our proposed CurNM method employs a dynamically updated negative pool and a curriculum learning strategy to adaptively adjust the difficulty level of negative samples during training, thus addressing the common challenge of positive sparsity. Furthermore, we implement a temporal-aware negative selection function to effectively capture positive shift. Our method’s efficacy is demonstrated through extensive experiments against 3 algorithms and 12 datasets. Although CurNM demonstrates superior performance, it requires tuning several hyperparameters to optimize its performance. In the future, we aim to simplify these designs to reduce the need for extensive parameter tuning.
@@ Acknowledgements
This work was supported by the Zhejiang Provincial Natural Science Foundation of China [Grant No. LMS25F020012], the Joint Funds of the Zhejiang Provincial Natural Science Foundation of China [Grant No. LHZSD24F020001], National Natural Science Foundation of China [Grant No. 62476238], Zhejiang Province High-Level Talents Special Support Program “Leading Talent of Technological Innovation of Ten-Thousands Talents Program” [Grant No. 2022R52046], and the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.